OpenStudy (anonymous):

simplify: (-4i)^5 ....why isn't the answer 16i

OpenStudy (mertsj):

Because (-4)^5=-1024

OpenStudy (mertsj):

and i^5=i

OpenStudy (anonymous):

hmm interesting okay, thanks!

OpenStudy (anonymous):

What part is interesting? 4^2 is 16. 4^5 = 4*4*4*4*4

OpenStudy (anonymous):

i did.. i^4-i=(i^2)^2 =(-4(-1))^2-i =16i

myininaya (myininaya):

i'm sorry alison I don't understand that type of reasoning \[(-4i)^5=(-4)^5(i)^5=(-4)(-4)(-4)(-4)(-4)i^{4+1}=-1024i^4i=-1024(1)i=-1024i\]

myininaya (myininaya):

my answer got cutoff -1024i

OpenStudy (anonymous):

Thanks okay, I understand your way better!